OpenStudy (anonymous):

What is the perimeter of the rectangle? Use the distributive property to rewrite the expression and solve. p=2l+2w =2*4.1+2*10.5 =2*(4.1+_____) = ______ft

OpenStudy (anonymous):

p=2l+2w = 2*4.1+2*10.5 = 2*(4.1+10.5) = 29.2